OpenStudy (anonymous):

how do you solve for x in \frac{1}{2(x-1)}=5+\frac{3}{x-1}

OpenStudy (lgbasallote):

\[\frac 12 (x - 2) = 5 + \frac 3{x-1}\] is that the question?

OpenStudy (anonymous):

yes that is the question

OpenStudy (lgbasallote):

okay...first multiply all terms by 2..what do you get?

OpenStudy (anonymous):

\[\frac{ 2 }{ 2(x-1)? }=10+\frac{6}{2(x-1)}

OpenStudy (lgbasallote):

yes that is right... except for one minor mistake \[\frac 22(x-2) = 10 + \frac{6}{x-1}\] do you see where you went wrong?

OpenStudy (anonymous):

i didn't multiply the negative 1 by 2. wouldn't i also multiply the x-1 by 2 as well

OpenStudy (lgbasallote):

no...remember i said multiply 2 to every term. 3/(x-1) is just one term so you multiply 2 to it only once. so you have \[\implies \frac 3{x-1} \times 2 \implies \frac 6{x-1}\] make sense?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

wonderful. so do you now get how i got \[\frac 22 (x-2) = 10 + \frac 6{x-1}\]

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

good. you can still simplify 2/2

OpenStudy (anonymous):

which would equal 1